Exercise 10-5 Working Backwards from Labor Variances [LO10-2]

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 2.50 $34.00 $85.00

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 54 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 350 F
Labor spending variance $ 500 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1
Labor efficiency variance = Labor spending variance +Labor rate variance = 500+350 = $850 U
1
Labor efficiency variance = SR*(AH-SH)
850 = 34.00*(AH-54*2.50)
25 = AH-135
AH = 135+25 = 160
Actual labor-hours = 160 hours
2
Labor rate variance = AH*(SR-AR)
350 = 160*(34.00-AR)
2.19 = 34.00-AR
AR = 34.00-2.19= $31.81
Actual hourly rate = $31.81